By a reversible adiabatic expansion, temperature of an ideal gas (n mol) changed from T, to Ts Express the changes of the internal energy and enthalpy by using heat capacity at conatant volume, Cy 2.
Answer #1

Change in internal energy, ΔU = nCv (T2-T1)

where n is the number of moles of ideal gas, T1 and T2 are the initial and final temperatures.

Change in enthalpy, ΔH = nCp (T2-T1) = n(Cv +R) (T2-T1)